# Structural barriers to complete homogenization and wormholing in dissolving porous and fractured rocks

### v1

@@ Abstract
Dissolution in porous media and fractured rocks alters both the chemical composition of the fluid and the physical properties of the solid. Depending on system conditions, reactive flow may enlarge pores uniformly, widen pre-existing channels, or trigger instabilities that form wormholes. The resulting pattern reflects feedbacks among advection, diffusion, surface reaction, and the initial heterogeneity of the medium. Porous and fractured media can exhibit distinct characteristics—for example, the presence of large fractures can significantly alter the network topology and overall connectivity of the system. We quantify these differences with three network models—a regular pore network, a disordered pore network, and a discrete fracture network—evaluated with a unified metric: the flow focusing profile. This metric effectively captures evolution of flow paths across all systems: it reveals a focusing front that propagates from the inlet in the wormholing regime, a system-wide decrease in focusing during uniform dissolution, and the progressive enlargement of pre-existing flow paths during channelization. The metric shows that uniform dissolution cannot eliminate heterogeneity resulting from the network topology. This structural heterogeneity—rather than just pore-diameter or fracture-aperture variance–sets a fundamental limit on flow homogenization and must be accounted for when upscaling dissolution kinetics from pore or fracture scale to the reservoir level.
@@
journal: International Journal of Rock Mechanics and Mining Sciences
@@
[C] Preprint Published in International Journal of Rock Mechanics and Mining Sciences (2026). DOI: 10.1016/j.ijrmms.2026.106431
@@ 1 Introduction
The flow of reactive fluids through rock induces significant changes in both the fluid composition and the properties of the solid matrix, which affects the transport dynamics within the system. Understanding these processes is essential for a wide range of geological and environmental applications. For instance, the safe disposal of high-level radioactive waste depends on the ability to predict the dominant modes of transport in the reservoir [76, 11]. Likewise, assessing contaminant migration requires understanding how flow is focused, because residence time determines whether a pollutant can be neutralized [24, 45, 52]. In both contexts, transport is tightly coupled to the evolving pore structure: reactive fluid remodels the medium and creates preferential pathways that ultimately control system behavior [5, 44, 33]. In the context of $\textrm{CO}_{2}$ sequestration, understanding both micro- and macro-scale behavior is critical for process optimization [3, 51, 69, 36, 58]. Predicting the evolution of flow pathways enables the avoidance of clogging and maximization of mineral replacement [7]. It is therefore evident that elucidating the coupling between flow, transport, and the evolving properties of the reservoir is essential to explain the diversity of structures that emerge from dissolution [9, 14, 32, 27].
@@
A wide body of experimental and numerical work has quantified how dissolution patterns depend on the relative time scales of advection, diffusion, and reaction. In porous media, dissolution experiments have investigated the pore scale behavior [57, 55, 22] and patterns emerging at the core scale [32, 27, 23, 26, 60, 67, 13]. Their findings have been reproduced and extended by multi-scale simulations, from direct numerical models to continuum codes [32, 57, 62, 74, 30, 68, 6, 56]. Four principal dissolution regimes are now recognized: compact, wormholing, channeling, and uniform. In compact (face) dissolution, the reactant is exhausted almost immediately after entering the sample, and a reaction front advances uniformly across its width [32, 12]. When wormholing occurs, advective instabilities localize flow into a few highly conductive channels that propagate from inlet to outlet [74, 59, 53]. In uniform dissolution, the reactant penetrates the full length of the system, enlarging pores homogeneously [27, 63]. Finally, in the channeling regime, initially existing flow pathways are uniformly enlarged along their entire length [56, 54]. Because the transitions between these regimes are subtle, distinguishing between them requires quantitative analysis of the system evolution [39, 70].
@@
Analogous dissolution phenomena also occur in fractured media, where high-permeability discontinuities channel most of the flow, distinguishing them from conventional porous media. Studies of dissolution in single-fracture systems—using both numerical models and flow-through experiments—have extensively documented aperture growth and channel development. [26, 19, 20, 21, 16, 46, 73].Similarly to porous media, studies of two-dimensional fracture planes have revealed compact, wormhole, and uniform dissolution regimes, each governed by the interplay of flow, transport, and chemical reactions [74, 29, 17]. Although the coupled flow–transport–reaction equations can be solved directly for idealized one- or two-dimensional fractures, doing so for natural systems with thousands of intersecting fractures—such as karst conduit networks—remains computationally formidable. To address this challenge, researchers employ a hierarchy of different representations: two-dimensional grids of intersecting fractures [20, 25, 18], fully three-dimensional networks with spatially variable apertures [47, 40, 41], and stochastic discrete-fracture-network (DFN) models [8, 37]. DFNs are attractive because they incorporate field-derived statistics of fracture aperture, length, and orientation, bridging single-fracture physics and network-scale heterogeneity while remaining computationally tractable for reactive-transport simulations [61].
@@
Porous media and fractured media exhibit fundamentally different physical structures and flow behaviors—fractures, for instance, can generate long-range, highly connected flow pathways—yet the impact of these distinct forms of heterogeneity on dissolution dynamics remains poorly understood and has not been systematically compared. Early work on variable-aperture fractures showed that the volume of reactant required for breakthrough varies non-monotonically with the magnitude of the initial aperture variation; a critical level of heterogeneity minimizes the breakthrough volume [29, 10]. The same trend was reproduced in numerical analysis of porous media that imposed spatially correlated porosity fields [42, 49, 50]. Follow-up studies examined the role of heterogeneity length scale [50] and suggested that the strong heterogeneity–wormholing link observed in small domains diminishes in larger systems [77]. There is, however, agreement that heterogeneity controls competition among channels and promotes tip branching during wormhole growth [10, 77].
@@
The interplay between heterogeneity and evolving dissolution patterns strongly affects the hydraulic properties of a medium. Pore-network simulations show that uniform dissolution can eliminate all variability in pore diameters [63, 15], whereas dissolving an initially homogeneous medium can instead generate pronounced heterogeneity; these opposing outcomes, in turn, remove or create anomalous transport behavior [5, 15, 64]. Heterogeneity in porous media has also been shown to affect the onset of preferential flow [13, 54, 1], permeability evolution [48], mixing [65], and even the overall dissolution regime [56, 70, 43]. These results show that significant changes in system dynamics can arise solely from adjusting pore-diameter variability. Beyond pore-size variability, heterogeneity also stems from the topology and geometry of pore or fracture networks—specifically, differences in connectivity and in segment-length distributions. This structural heterogeneity also shapes system evolution, influencing processes ranging from karst genesis [2] to mineral carbonation [38].
@@
In this paper, we quantify how different forms of heterogeneity couple with dissolution across a range of flow and reaction conditions. Three network models are analyzed: a regular pore network with variance in pore diameters, a disordered pore network with variance in pore diameters and lengths, and a discrete fracture network with variance in fracture apertures, lengths, and connectivity. We distinguish three heterogeneity types: conduit-scale (pore diameter or fracture aperture), segment-scale (length), and network-scale (connectivity). Tracking each scale through time shows that initial heterogeneity shapes the evolution of flow pathways and ultimately limits the degree to which dissolution can homogenize the system.
@@ 2 Reactive Transport in Network Models
We quantitatively describe the dissolution regimes and analyze their transport properties by utilizing network models. We use a capillary pore network model for the dissolving porous medium [32, 6] and a graph representation of a discrete fracture network for dissolving fractures [8, 37]. We adjust the initial properties, that is, connectivity, lengths of connections, and distributions of pore diameters and fracture apertures, to exhibit the differences in dissolution arising from the structure of the medium.
@@ 2.1 Network Generation
    Regular pore network – a diamond lattice in which every connection has identical length; variability arises solely from the log-normal distribution of pore diameters (conduit-scale heterogeneity).
@@
    Disordered pore network – a Delaunay network whose nodes are randomly positioned, introducing heterogeneity in pore lengths and intersection angles while retaining the same diameter distribution as the regular lattice (conduit- and segment-scale heterogeneity).
@@
    Discrete fracture network – a semi-generic model loosely based on the fractured carbonate-hosted Pietrasecca Fault in the central Apennines, Italy [66]; node locations, fracture intensities, and aperture statistics reproduce the multiscale heterogeneity of natural carbonate fractures (conduit-, segment-, and network-scale heterogeneity).
Figure 1 shows a representative example of each network.
Figure 1: (a) Example regular pore network realization (diamond lattice); edge width is proportional to initial pore diameter. (b) Example disordered pore network realization (Delaunay lattice); edge width again reflects initial pore diameter. (c) Example discrete fracture network; colors denote fracture families. (d) Graph representation of the same DFN, projected into two dimensions via principal-component analysis.
For every geometry we generate 30 statistically equivalent realizations. In the two pore network cases we use a domain of side length $L$ discretized into $100\times 100$ nodes; periodic boundaries are applied perpendicular to the imposed pressure gradient. For the regular pore network, each realization shares the same connectivity but draws pore diameters from an independent log-normal distribution with mean $d_{0}$, variance $d_{0}^{2}$ (that is, the coefficient of variation is $\sim 1$), and correlation length of $0.1L$; the sampling procedure follows [70, 77]. For the disordered pore network, the nodes are randomly distributed on the domain and connected using Delaunay triangulation; node positions are different for each realization. Pore diameters are assigned from the same distributions as for the regular network.
The DFN domain (25 m × 10 m × 10 m) is populated with three fracture families that differ in orientation, length distribution, aperture statistics, and intensity; full generation parameters are listed in A. Importantly, each fracture family has a set aperture, constant throughout the realizations, but the proportions between families vary. We convert each DFN to a pipe-network graph with the algorithm of [35], preserving fracture intersections as graph nodes.
In every model the nodes on one boundary serve as the inlet, while those on the opposite boundary form the outlet. In the pore network cases we impose periodic boundary conditions in the direction perpendicular to the inlet–outlet axis.
@@ 2.2 Reactive Transport
Our reactive transport model follows the derivations of [6] for pore networks and [75] for single fractures. Volumetric flow $q$ in each channel obeys
@@
where $C$ is the hydraulic conductance, $P$ the pressure, and $\mu$ the fluid viscosity. For a cylindrical pore $C=\pi d^{4}/128$, with $d$ the pore diameter (Hagen–Poiseuille equation); for a fracture slot $C=wb^{3}/12$, with $w$ the fracture width and $b$ the aperture (Reynolds equation). A constant total volumetric flow rate $Q$ is enforced by setting $P=P_{0}$ at the inlet nodes and $P=0$ at the outlet, then rescaling $P_{0}$ at every time step to maintain $Q$.
A reactant of concentration $c_{\textrm{in}}$ is injected into the inlet nodes. Along each connection we solve the one-dimensional advection–reaction equation
$q\frac{dc}{dx}=-sk_{\textrm{eff}}c,$ (2)
where $x$ denotes the direction along the channel length, the reactive surface is $s=\pi d$ for pores and $s=2w$ for fractures, and $k_{\textrm{eff}}$ is the effective reaction rate,
$k_{\textrm{eff}}=k/(1+kd^{\textrm{h}}/D\textrm{Sh}),$ (3)
which takes into account the hindering effects of diffusion across the pore diameter or fracture aperture. In Eq. (3), $k$ denotes the reaction rate, $D$ is the diffusion coefficient, and Sh is the Sherwood number [4, 31, 28]. Finally, $d^{\textrm{h}}$ is the hydraulic diameter, equal to $d$ for pores and $2b$ for fractures. The Sherwood number, Sh, depends on reaction rate at mineral surfaces, but the variation is relatively small, for porous media limited by the values $\textrm{Sh}=3.656$ and $\textrm{Sh}=4.364$ (we approximate it by a constant value, $\textrm{Sh}=4$), while in fractured media, from $\textrm{Sh}=7.54$ to $\textrm{Sh}=8.24$ (we approximate it by $\textrm{Sh}=8$).
@@
$\partial_{t}d=\frac{2k_{\textrm{eff}}(d)}{\nu c_{\textrm{sol}}}c,\ \ \partial_{t}b=\frac{2k_{\textrm{eff}}(b)}{\nu c_{\textrm{sol}}}c,$ (4)
where $c_{\textrm{sol}}$ is the concentration of soluble material and $\nu$ accounts for the stoichiometry of the reaction.

### v2

@@ Abstract
Dissolution in porous media and fractured rocks alters both the chemical composition of the fluid and the physical properties of the solid. Depending on system conditions, reactive flow may enlarge pores uniformly, widen pre-existing channels, or trigger instabilities that form wormholes. The resulting pattern reflects feedbacks among advection, diffusion, surface reaction, and the initial heterogeneity of the medium. Porous and fractured media can exhibit distinct characteristics—for example, the presence of large fractures can significantly alter the network topology and overall connectivity of the system. We quantify these differences with three network models—a regular pore network, a disordered pore network, and a discrete fracture network—evaluated with a unified metric: the flow focusing profile. This metric effectively captures evolution of flow paths across all systems: it reveals a focusing front that propagates from the inlet in the wormholing regime, a system-wide decrease in focusing during uniform dissolution, and the progressive enlargement of pre-existing flow paths in the channeling regime. The metric shows that uniform dissolution cannot eliminate heterogeneity resulting from the network topology. This structural heterogeneity—rather than just pore-diameter or fracture-aperture variance—sets a fundamental limit on flow homogenization and must be accounted for when upscaling dissolution kinetics from pore or fracture scale to the reservoir level.
@@
journal: IJRMMS
@@
[C] Accepted manuscript Published in International Journal of Rock Mechanics and Mining Sciences (2026). DOI: 10.1016/j.ijrmms.2026.106431 Licensed under CC BY-NC-ND 4.0 (https://creativecommons.org/licenses/by-nc-nd/4.0/).
@@ 1 Introduction
The flow of reactive fluids through rock induces significant changes in both the fluid composition and the properties of the solid matrix, thereby affecting the transport dynamics within the system. Understanding these processes is essential for a wide range of geological and environmental applications. For instance, the safe disposal of high-level radioactive waste depends on the ability to predict the dominant modes of transport in the reservoir. Tsang et al. [2015], Churakov et al. [2020] Likewise, assessing contaminant migration requires understanding how flow is focused, because residence time determines whether a pollutant can be neutralized. Fryar and Schwartz [1998], Kung et al. [2000], Mayer et al. [2001] In both contexts, transport is tightly coupled to the evolving pore structure: reactive fluid remodels the medium and creates preferential paths that ultimately control system behavior. Berkowitz and Scher [1998], Kang et al. [2015], Hyman et al. [2016] In the context of $\textrm{CO}_{2}$ sequestration, understanding both micro- and macro-scale behavior is critical for process optimization. Andreani et al. [2009], Matter and Kelemen [2009], Steefel et al. [2013], Hyman et al. [2020], Nisbet et al. [2024] Predicting the evolution of flow paths enables the avoidance of clogging and maximization of mineral replacement. Budek and Szymczak [2025] It is therefore evident that elucidating the coupling between flow, transport, and the evolving properties of the reservoir is essential to explain the diversity of structures that emerge from dissolution. Chadam et al. [1986], Daccord and Lenormand [1987], Hoefner and Fogler [1988], Golfier et al. [2002]
@@
A wide body of experimental and numerical work has quantified how dissolution patterns depend on the relative time scales of advection, diffusion, and reaction. In porous media, dissolution experiments have investigated the pore scale behavior Molins et al. [2014], Menke et al. [2015], Dutka et al. [2020] and patterns emerging at the core scale. Hoefner and Fogler [1988], Golfier et al. [2002], Fredd and Fogler [1998], Garcia-Rios et al. [2017], Ott and Oedai [2015], Snippe et al. [2020], Cooper et al. [2023] Their findings have been reproduced and extended by multi-scale simulations, from direct numerical models to continuum codes. Hoefner and Fogler [1988], Molins et al. [2014], Panga et al. [2005], Szymczak and Ladd [2009], Hao et al. [2013], Steefel et al. [2015], Budek and Szymczak [2012], Menke et al. [2023] Four principal dissolution regimes are now recognized: compact, wormholing, channeling, and uniform. In compact (face) dissolution, the reactant is exhausted almost immediately after entering the sample, and a reaction front advances uniformly across its width. Hoefner and Fogler [1988], Cohen et al. [2008] When wormholing occurs, advective instabilities localize flow into a few highly conductive channels that propagate from inlet to outlet. Szymczak and Ladd [2009], Ortoleva et al. [1987], McDuff et al. [2010] In uniform dissolution, the reactant penetrates the full length of the system, enlarging pores homogeneously. Golfier et al. [2002], Roded et al. [2020] Finally, in the channeling regime, initially existing flow paths are uniformly enlarged along their entire length. Menke et al. [2023, 2016] Because the transitions between these regimes are subtle, distinguishing between them requires quantitative analysis of the system evolution. Jang et al. [2011], Szawełło et al. [2024]
@@
Analogous dissolution phenomena also occur in fractured media, where high-permeability discontinuities channel most of the flow, distinguishing them from conventional porous media. Studies of dissolution in single-fracture systems—using both numerical models and flow-through experiments—have extensively documented aperture growth and channel development. Garcia-Rios et al. [2017], Dreybrodt [1990, 1996], Durham et al. [2001], Detwiler et al. [2003], Li et al. [2008], Szymczak and Ladd [2004] Similarly to porous media, studies of two-dimensional fracture planes have revealed compact, wormhole, and uniform dissolution regimes, each governed by the interplay of flow, transport, and chemical reactions. Szymczak and Ladd [2009], Hanna and Rajaram [1998], Detwiler and Rajaram [2007] Although the coupled flow–transport–reaction equations can be solved directly for idealized one- or two-dimensional fractures, doing so for natural systems with thousands of intersecting fractures—such as karst conduit networks—remains computationally formidable. To address this challenge, researchers employ a hierarchy of different representations: two-dimensional grids of intersecting fractures, Dreybrodt [1996], Gabrovšek and Dreybrodt [2001], Dreybrodt et al. [2005] fully three-dimensional networks with spatially variable apertures, Li et al. [2020], Jiang et al. [2022, 2023] and stochastic discrete fracture network (DFN) models. Cacas et al. [1990], Hyman et al. [2015] DFNs are attractive because they incorporate field-derived statistics of fracture aperture, length, and orientation, bridging single-fracture physics and network-scale heterogeneity while remaining computationally tractable for reactive transport simulations. Pachalieva et al. [2025]
@@
Porous media and fractured media exhibit fundamentally different physical structures and flow behaviors—fractures, for instance, can generate long-range, highly connected flow paths—yet the impact of these distinct forms of heterogeneity on dissolution dynamics remains poorly understood and has not been systematically compared. Early work on variable-aperture fractures showed that the volume of reactant required for breakthrough varies non-monotonically with the magnitude of the initial aperture variation; a critical level of heterogeneity minimizes the breakthrough volume. Hanna and Rajaram [1998], Cheung and Rajaram [2002] The same trend was reproduced in numerical analysis of porous media that imposed spatially correlated porosity fields. Kalia and Balakotaiah [2009], Maheshwari et al. [2013], Maheshwari and Balakotaiah [2013] Follow-up studies examined the role of heterogeneity length scale Maheshwari and Balakotaiah [2013] and suggested that the strong heterogeneity–wormholing link observed in small domains diminishes in larger systems. Upadhyay et al. [2015] There is, however, agreement that heterogeneity controls competition among channels and promotes tip branching during wormhole growth. Cheung and Rajaram [2002], Upadhyay et al. [2015]
@@
The interplay between heterogeneity and evolving dissolution patterns strongly affects the hydraulic properties of a medium. Pore network simulations show that uniform dissolution can eliminate all variability in pore diameters, Roded et al. [2020], Deng et al. [2025] whereas dissolving an initially homogeneous medium can instead generate pronounced heterogeneity; these opposing outcomes, in turn, remove or create anomalous transport behavior. Berkowitz and Scher [1998], Deng et al. [2025], Saeibehrouzi et al. [2025] Heterogeneity in porous media has also been shown to affect the onset of preferential flow, Cooper et al. [2023], Menke et al. [2016], Al-Khulaifi et al. [2019] permeability evolution, Li et al. [2019] mixing, Sharma et al. [2023] and even the overall dissolution regime. Menke et al. [2023], Szawełło et al. [2024], Kanavas et al. [2025] These results show that significant changes in system dynamics can arise solely from adjusting pore-diameter variability. Beyond pore-size variability, heterogeneity also stems from the topology and geometry of pore or fracture networks—specifically, differences in connectivity and in segment-length distributions. This structural heterogeneity also shapes system evolution, influencing processes ranging from karst genesis Aliouache and Jourde [2024] to mineral carbonation. Hyman et al. [2024]
@@
In this paper, we utilize the flow focusing metric introduced by Szawełło et al. Szawełło et al. [2024] to quantify how different forms of heterogeneity couple with dissolution across a range of flow and reaction conditions. We apply the metric to three network models: a regular pore network with variance in pore diameters, a disordered pore network with variance in pore diameters and lengths, and a discrete fracture network with variance in fracture apertures, lengths, and connectivity. We distinguish three heterogeneity types: conduit-scale (pore diameter or fracture aperture), segment-scale (length), and network-scale (connectivity). By tracking each scale through time, we show that initial heterogeneity shapes the evolution of flow paths and ultimately limits the degree to which dissolution can homogenize the system. Our focus is on the impact of length and connectivity, which together form a quenched disorder in the network—that is, a structural heterogeneity that dissolution can modify only marginally. Initially, this disorder plays a smaller role than the widely studied heterogeneity in conduit width: pore diameters enter the conductance in the fourth power and fracture apertures in the third, whereas lengths enter only linearly. As homogenization progresses and variation in conduit widths diminishes, however, segment- and network-scale heterogeneity become the decisive factors in flow-path selection. This behavior is key to understanding the differing responses of porous media and fracture networks to dissolution.
@@ 2 Reactive Transport in Network Models
We quantitatively describe the dissolution regimes and analyze their transport properties by utilizing network models. We use a capillary pore network model for the dissolving porous medium Hoefner and Fogler [1988], Budek and Szymczak [2012] and a graph representation of a discrete fracture network for dissolving fractures. Cacas et al. [1990], Hyman et al. [2015] We adjust the initial properties, that is, connectivity, conduit lengths, and distributions of pore diameters and fracture apertures, to exhibit the differences in dissolution arising from the structure of the medium.
@@ 2.1 Network Generation
    Regular pore network: a diamond lattice in which every edge has identical length; variability arises solely from the log-normal distribution of pore diameters (conduit-scale heterogeneity).
@@
    Disordered pore network: a Delaunay network whose nodes are randomly positioned, introducing heterogeneity in pore lengths and intersection angles while retaining the same diameter distribution as the regular lattice (conduit- and segment-scale heterogeneity).
@@
    Discrete fracture network: a semi-generic model loosely based on the fractured carbonate-hosted Pietrasecca Fault in the central Apennines, Italy Smeraglia et al. [2021]; node locations, fracture intensities, and aperture statistics reproduce the multiscale heterogeneity of natural carbonate fractures (conduit-, segment-, and network-scale heterogeneity).
Fig. 1 shows a representative example of each network.
Figure 1: (a) Example regular pore network realization (diamond lattice); edge width is proportional to the initial pore diameter. (b) Example disordered pore network realization (Delaunay lattice); edge width again reflects the initial pore diameter. (c) Example discrete fracture network; colors denote fracture families. (d) Graph representation of the same DFN, projected into two dimensions via principal component analysis. Edge width is kept constant, so color intensity corresponds to the local density of fracture segments.
For every geometry, we generate 30 statistically equivalent realizations. In the two pore network cases, we use a domain of side length $L$ discretized into $100\times 100$ nodes. Periodic boundary conditions are applied in the direction perpendicular to the imposed pressure gradient: pores on one side of the lattice are connected to their counterparts on the opposite side, which removes artificial side-wall effects.
For the regular pore network, each realization shares the same connectivity, but draws pore diameters from an independent log-normal distribution with mean $d_{0}$, variance $d_{0}^{2}$ (that is, the coefficient of variation is $\sim 1$), and correlation length of $0.1L$; the sampling procedure follows Szawełło et al. Szawełło et al. [2024] and Upadhyay et al. Upadhyay et al. [2015] The value of $d_{0}$ can range from micrometers to millimeters, depending on the mineral and rock type; because the model is formulated in dimensionless variables, its exact value is not prescribed. The formulation is, however, most appropriate for long, slender pores, so that dissolution does not drastically alter the network topology. In the regular network, all pore lengths are set equal to a reference length $l_{0}$.
For the disordered pore network, nodes are randomly distributed in the domain and connected using a Delaunay triangulation; node positions differ between realizations. To impose periodic boundary conditions, we tile the original domain with its translated copies and construct a Delaunay triangulation on this enlarged set of points. From the resulting edges, we keep only those whose midpoints lie inside the original domain. For edges that cross a side boundary, we connect them to the matching points on the opposite side, creating periodic links. Pore diameters are assigned from the same distributions as for the regular network, while pore lengths are given by the Euclidean distance between connected nodes; by construction, the average pore length is equal to $l_{0}$.
The DFN domain (25 m × 10 m × 10 m) is populated with three fracture families that differ in orientation, length distribution, aperture statistics, and intensity; full generation parameters are listed in A. Importantly, each fracture family has a set aperture, constant throughout the realizations, but the proportions between families vary. We convert each DFN to a pipe-network graph with the algorithm of Hyman et al., Hyman et al. [2018] preserving fracture intersections as graph nodes.
In every model, the nodes on one boundary serve as the inlet, while those on the opposite boundary form the outlet.
@@ 2.2 Reactive Transport
Our reactive transport model follows the derivations of Budek and Szymczak Budek and Szymczak [2012] for pore networks and Szymczak and Ladd Szymczak and Ladd [2012] for single fractures. In the pore network representations (Fig. 1a,b), each edge corresponds to a cylindrical pore of diameter $d$ and length $l$; in the discrete fracture network (Fig. 1c,d), each edge represents a fracture segment of aperture $b$ and length $l$ spanning a fracture of width $w$. Volumetric flow $q$ in each channel obeys
@@
where $C$ is the hydraulic conductance, $P$ the pressure, and $\mu$ the fluid viscosity. For a cylindrical pore $C=\pi d^{4}/128$ (Hagen–Poiseuille equation); for a fracture slot $C=wb^{3}/12$ (Reynolds equation). At each intersection $i$, we impose mass conservation in the form of the nodal continuity condition
$\sum_{j}q_{ij}=0,$ (2)
where the sum is over all edges $ij$ incident on node $i$. Across the network, a constant total volumetric flow rate $Q$ is enforced by setting $P=P_{0}$ at the inlet nodes and $P=0$ at the outlet, then rescaling $P_{0}$ at every time step to maintain $Q$.
We consider dissolution of a porous medium or a fracture network by an aqueous reactant of inlet concentration $c_{\textrm{in}}$, injected at the inlet nodes. We assume a single-component dissolution reaction with a linear rate law
$R(c_{\textrm{w}})=kc_{\textrm{w}},$ (3)
where $k$ is the surface reaction rate and $c_{\textrm{w}}$ is the reactant concentration at the pore/fracture wall. This rate law is a standard approximation for far-from-equilibrium reactions, such as dissolution by concentrated acid; in that case $c_{\textrm{w}}$ corresponds to the concentration of $\textrm{H}^{+}$ ions. Hoefner and Fogler [1988] An equivalent formulation appears in the karst literature, where limestone dissolution is often treated as first order in undersaturation, which maps directly onto our driving concentration variable. Dreybrodt [1990]
It is convenient to express the wall concentration in terms of the bulk concentration $c$. The rate law in Eq. (3) can then be written as $R(c)=k_{\textrm{eff}}(d^{\textrm{h}})c$, where $k_{\textrm{eff}}(d^{\textrm{h}})$ is an effective reaction rate that incorporates the hindering effects of transverse diffusion:
$k_{\textrm{eff}}(d^{\textrm{h}})=\frac{k}{1+\frac{kd^{\textrm{h}}}{D\textrm{Sh}}},$ (4)
where $D$ is the diffusion coefficient, Sh is the Sherwood number, Bejan [1984], Hayes and Kolaczkowski [1994], Gupta and Balakotaiah [2001] and $d^{\textrm{h}}$ is the hydraulic diameter, equal to $d$ for pores and $2b$ for fractures. The Sherwood number, Sh, depends on reaction rate at mineral surfaces, but the variation is relatively small, for porous media limited by the values $\textrm{Sh}=3.656$ and $\textrm{Sh}=4.364$ (we approximate it by a constant value, $\textrm{Sh}=4$), while in fractured media, from $\textrm{Sh}=7.54$ to $\textrm{Sh}=8.24$ (we approximate it by $\textrm{Sh}=8$).
Along each conduit we solve the one-dimensional advection–reaction equation for the bulk concentration
$q\frac{dc}{dx}=-sk_{\textrm{eff}}(d^{\textrm{h}})c,$ (5)
where $x$ is the coordinate along the channel length, the reactive surface is $s=\pi d$ for pores and $s=2w$ for fractures.
While in Eq. (5) the effective reaction rate $k_{\textrm{eff}}(d^{\textrm{h}})$ incorporates the effects of transverse diffusion, axial diffusion (along the flow direction) is neglected. In general, this approximation is justified when the Péclet number, which measures the relative importance of advection and diffusion, is larger than one. The definition of the Péclet number requires a choice of length scale; for reactive infiltration this scale is set by the penetration length of the reactant, that is, the characteristic distance it can travel into the medium before being consumed. If, on this length scale, advection dominates over axial diffusion, then axial diffusion can be safely neglected.
At each intersection $i$, we assume perfect mixing of the concentration,
$c_{i}=\frac{\sum_{j^{\prime}}q_{ij}c_{ij}^{\textrm{out}}}{\sum_{j^{\prime}}q_{ij}},$ (6)
where the sum runs over all edges with flow directed into node $i$ and $c_{ij}^{\textrm{out}}$ is the outlet concentration at the end of those edges. The node concentration $c_{i}$ is then used as the inlet concentration for all edges with flow directed out of node $i$.
@@
$\partial_{t}d=\frac{2k_{\textrm{eff}}(d)}{\nu c_{\textrm{sol}}}\,c,\qquad\partial_{t}b=\frac{2k_{\textrm{eff}}(2b)}{\nu c_{\textrm{sol}}}\,c,$ (7)
where $c_{\textrm{sol}}$ is the molar concentration of soluble mineral in the solid and $\nu$ is the stoichiometric coefficient of the reaction. To keep the model tractable, in each time step we compute the total volume of mineral dissolved along the conduit and increase its diameter (or aperture) uniformly so that the conduit volume increases by exactly that amount, thereby preserving mass conservation.
The model assumes a separation of time scales between transport relaxation and mineral dissolution: for each geometry we solve the steady-state flow and transport problem (Eq. (5)) and then update $d$ or $b$ using the above law. This quasi-steady approximation is appropriate when the acid capacity number $\gamma=c_{\textrm{in}}/(\nu c_{\textrm{sol}})$, which measures the volume of solid dissolved by a unit volume of reactant, is much smaller than one. This condition is satisfied in the great majority of geological and industrial reactive transport processes. There are, however, systems where this approximation can break down, for example in the dissolution of halite or caramel. Cohen et al. [2020]